what is impulse how is it different from momentum

Grade:12

1 Answers

Ramesh V
14 years ago

Momentum:

The momentum of a body is equal to its mass multiplied by its velocity.Momentum is measured in N-s. and momentum is a vector quantity.

Impulse:

The impulse of a force is equal to the change in momentum of a body which a force causes. This is also equal to the magnitude of the force multiplied by the length of time the force is applied.

                                                                  Impulse = change in momentum = force × time

These concepts are merely an outgrowth of Newton's second law  (Fnet = m • a),which stated that the acceleration of an object is directly proportional to the net force acting upon the object and inversely proportional to the mass of the object. When combined with the definition of acceleration (a = change in velocity / time), the following equalities result.

                                                                                    F=m*a = m*dV/t

If both sides of the above equation are multiplied by the quantity t, a new equation results.

                                                                                       F*t  = m*dV

This equation represents one of two primary principles to be used in the analysis of collisions during this unit. To truly understand the equation, it is important to understand its meaning in words. In words, it could be said that the force times the time equals the mass times the change in velocity. In physics, the quantity Force • time is known as impulse. And since the quantity m•v is the momentum, the quantity m•v must be the change in momentum.

The equation really says that the : Impulse = change in momentum

 

Example: To understand the physics of collisions.

In a collision, an object experiences a force for a specific amount of time which results in a change in momentum. The result of the force acting for the given amount of time is that the object's mass either speeds up or slows down (or changes direction). The impulse experienced by the object equals the change in momentum of the object.
In equation form, F • t = m • v.